Back to the NIKA2 main page

Daily Reports

Nov 8th, Tuesday

16:30-21:00; Ioannis & Alice: Maintenance finished at 16:30 but wind was too high to start with observations. Wind speed reduced a bit by 18:00 and started system check with EMIR because Tau225 was ~0.4-0.5 at that time and expected to increase further. Observations for EMIR backup project 056-22 between 19:00 and 21:00 for objects G053.23 and G49.48 until atmosphere became too unstable to continue (pwv fluctuations between 6mm and 10mm). All NIKA2 *.ini files and IMBFITS creation scripts were updated for ongoing run73.

21:00-03:00; Hélène: Tried a couple of pointings but atmosphere was too unstable. Observations stopped at 21:50 due to high wind speed.

Nov 9th, Wednesday

03:00-09:00; Philippe & Tuan-Anh: EMIR backup observations throughout the shift, first for project 051-22 (one set of maps on NGC1977) and then for project 079-22 (a couple of repetitions on P07-C and P01-C). Opacity high and quite variable (Tau225 ~ 0.5-0.8).

09:00-15:00; Juan Macías & Alejandro Jimenez: EMIR backup observations for project 083-22 (39573-high; 2 sets of repetitions) until 12:30. Stopped after that due to very unstable atmosphere and fog (prediction shows showers this afternoon).

15:00-21:00; Alice: Weather too unstable (Tau225 ~ 0.5 - 1) and a bit of snowfall. No observations.

21:00-03:00; Hélène & Juan: Opacity decreased and we switch finally to NIKA2! :-) Tested acquisition in both total power and polarization mode and system seems to work fine. Bad KIDs are about 70/5/90 for Ar1/Ar2/Ar3. After several tunings the bad KIDs of Ar3 remain at this relatively high level but not too much (80-90). Several pointing, focus and calibration scans on Uranus while we wait for the opacity to decrease further. Note that calib_1scan scans 20221109s150 and skydip scan 20221109s151 were taken at relatively high opacity and not too stable conditions and calib_1scan scans 20221109s161 and 20221109s162 were taken with the HWP in front of the cryostat window. Around 22:30 UT the opacity reaches a plateau around 0.3-0.35 and baseline variations are < 1 Jy/beam at 2mm, so we do a proper calib_1scan [20221109s171] and a skydip at lower opacity [20221109s172] and we start with observations for the GASTON LP (122-16).

Hélène's input: The observations actually started at ~23:45 local time, after a series of calibrations on Uranus, when the opacity decreased sufficiently. It remained around 0.3, with occasional big spikes. Did 2 repetitions on L1521E for project 122-16 at 60"/s, pointing and focus on Uranus at ~01:15 (the focus changed from -0.54 to -0.73), one scan on L1521E at 40"/s, after which the source was too high to continue. Pointing and focus again on Uranus at ~02:20 (the focus changed from -0.73 to -0.78) before doing a beammap [IM: scan 20221110s18]. Then one more scan on L1521E at 40"/s to get a complete repetition with both scan directions. The stability was difficult to gauge (observing remotely) and will have to be checked. The RTA was noticeably slower than in previous pools and this becomes a problem because some observing time is lost while waiting for the focus and pointing results. For some unknown reason, the 2mm data were not processed correctly by the pipeline, that was unable to provide pointing corrections. There were taken from the PIIC RTA.

Nov 10th, Thursday

03:00-09:00; Philippe & Tuan-Anh: We switched to NIKA2-pol around 03:30 and did a few calibration tests on Uranus and Mars (polarisation-mode calib_1scans : 20221110s[36,37,44,46]) and also a polarization mode skydip [20221110s45]. Focus was stable but opacity and atmospheric stability were marginal. nk_rta did not work in polarization mode (idl_polar). Then we switched to total power around 05:30 and after a couple of calib_1scans on Mars and CRL618 [20221110s53 and 20221110s54, respectively] we observed project 008-22 (a pair of maps on IC348-1, Tau225 ~ -0.2-0.25) and later project 095-22 (2 sequences on AR2623, Tau225~0.2).

09:00-15:00; Juan-Macías & Alex Jimenez: Continued with observations of project 095-22 and the target AR2623 until 10:00. Then moved to project 093-22 and observed PSZ2G229 for a couple of repetitions. A problem with the array2 electronics was identified but solved after rebooting the rack, we lost ~40 min. Around 12:35 the opacity started to increase (rising from Tau225 ~ 0.2 to 0.3) and we moved to project 032-22. After one repetition on J1509, the opacity increased further to > 0.4 and we switched to EMIR.

15:00-21:00; Alice & Juan: Weather deteriorated fast. Trying pointing and focus with EMIR 3mm band with no luck for the major part of the shift. Towards the end, opacity decreased and we did one repetition on G069.54 for the EMIR backup project 056-22.

21:00-03h00; Helene: During the run the taumeter opacity was around 0.30 but with some spiky behavior from time to time. The telescope was fine. Run started with pointing, focus, Calib_1scan [20221110s187] on URANUS, and skydip [20221110s188]. Then : (1) project 122-16 L1512E two complete observations (2 directions) at 60"/sec while source below 70deg, (2) pointing CRL618, (3) project 122-16 L1512E two complete observations (2 directions) at 60"/sec while source below 70deg, (4) pointing CRL618 (pointing was stable), (5) project 122-16 L1512E one observation in FIRST direction at 40"/sec while source below 75deg, (6) L1512E too high : so pointing, focus, Calib_1scan [20221111s16] on URANUS, and skydip [20221111s17] (focus changed from -0.33 to -0.45), (7) At this point I could not catch L1512E after transit for another observation in the SECOND direction before the end of the shift. (I should have made a beammap instead of attempting this tight sequence really).

Nov 11th, Friday

03:00-09:00; Philippe & Tuan-Anh: Atmospheric opacity was marginally OK for NIKA2, tau225 ~ 0.2-0.5 according to radiometer but with strong fluctuations likely due to a high level of local relative humidity (~ 75%) close to the telescope. Real opacity probably ~ 0.3. However, sky was clear with stable conditions (rms fluctuations between ~1 and ~3 Jy/beam at 2mm). We switched from NIKA2 TP to Polarimetry mode at ~ 3am (LT). Performed polarized calibration scans (calib_1scan) on Uranus (s27, s34, s40) and Mars (s28, s41, s48) and did a polarized skydip (s43). Took 1 sequence of 3 polarized maps on OMC1 (s30, s31, s32) and 1 sequence of 2 maps on OMC3 (s45, s46) for B-FUN (015-17). Then switched back to NIKA2 TP and observed 1 sequence of 11 maps on AR2623 for 095-22. Had to stop at ~ 8:15 am (LT) due to strong wind.

09:00-15:00; Juan-Macías & Alex Jimenez: Strong wind, telescope stopped.

15:00-21:00; Alice: Strong wind, telescope stopped.

21:00-03:00; Aurore: Strong wind, telescope stopped.

Nov 12th, Saturday

03:00-09:00; Philippe & Tuan-Anh: Strong wind and snowstorm, telescope stopped.

09:00-15:00; Juan-Macías: Strong wind and snowstorm, telescope stopped.

15:00-21:00; Alice: Strong wind, telescope stopped.

21:00-03:00; Aurore: Wind dropped below the limit close to the end of the shift. We started NIKA2, atmospheric stability was low until the end of the shift (+/- 8-10 Jy/beam @ 2mm) but opacity was low (Tau225 ~ 0.2), so we did some calibrations on Uranus: calib_1scan [20221113s16], beam map [20221113s18] and skydip [20221113s20].

Nov 13th, Sunday

03:00-09:00; Philippe & Tuan-Anh: Starting with NIKA2-POL at 4 am (LT) after a long period of strong wind and unstable weather. Clear sky, low atmospheric opacity tau225 ~ 0.2 or less. Stable conditions, getting slightly better with time: rms baseline fluctuations between ~1 Jy/beam (at start of shift) and ~0.5 Jy/beam (at end of shift) in pointing scans at 2mm. Performed polarized calibration scans (calib_1scan) on Uranus (s27) and Mars (s33, s49, s56) and did a polarized skydip (s66). Took 2 sequences of 2 polarized maps on OMC3 (s35, s36, s37, s38), 1 sequence of 3 polarized maps on OMC2 (s40, s41, s42), and 2 sequences of 3 + 2 polarized maps on NGC2264 for B-FUN (015-17). Switched to NIKA2 TP at ~8:30 am (no more B-FUN sources).

09:00-15:00; Juan-Macías & Alice: We started NIKA TP at around 8h45. After pointing and focus we moved to N2CLS and observed COSMOS for about 1h30. Weather conditions were reasonably good with tau < 0.25 although some instabilities. When COSMOS was low we moved to the 093-22 project and observed PSZ2G107 for two hours. The weather reasonable at the beginning but it degrades with time. Mainly large atmospheric fluctuations that showed up in the PIIC quick look maps. With the AOD we decided to move to project 032-22 and observed J1509 for about one hour. At the end of the shift we started to see issues with the beams. We observed very large beams of about 20 arsecs at 2 mm.

15:00-21:00; Alice: We tried to redo focus and pointing but beams too large as before. We switched to EMIR until the beam conditions improve and observed projects 019-22 and 056-22. At 20:00 we switch back to NIKA2 and verified that the beam are more symmetric and focus is stable. Then we started observations for project 008-22.

21:00-03:00; Jean-François: Shift started with a couple of maps for NIKA2 project 008-22 with opacity around 0.4. Then, flux calibration : pointing, focus, Calib_1scan [20221113s229] on URANUS, skydip [20221113s221]. At that point, opacity increased significantly, and we turn to EMIR 3mm. We did 6 maps for EMIR project 051-22 on source NGC1977 (about 2 hours). Telescope was fine.

Nov 14th, Monday

03:00-09:00; Philippe & Tuan-Anh: Observing with NIKA2 (TP). Clear sky, relatively high atmospheric opacity tau225 ~ 0.3-0.4, fairly stable conditions with rms baseline fluctuations ~ 0.5 to ~ 2.5 Jy/beam in pointing scans at 2mm. Performed calibration scans (calib_1scan) on Uranus (s44), Mars (s62), AlphaOri (s63), and did a skydip (s78). Executed 4 repetitions of the map.pako script (i.e., 8 scans) on source IC348-1 for project 008-22. Integrated for ~ 1 hr (1 repetition otfmap_J0842-0157.pako, 12 scans) on source J0842-0157 for project 089-22. Switched to project 032-22, observing source J1522, with Juan taking over at ~8:30 am (LT).

09:00-15:00; Juan & Alice: I took over the shift around 10 am. Begin with NIKA2 observation project 032-22. We observed J1510 for 4 rep and J1228 for 2 reps. Because of the unstable atmospheric conditions, we then switched to EMIR and observed project 083-22 and then 019-22 until the end of the shift.

15:00-21:00; Ioannis: Continued with EMIR backup projects due to high atmospheric fluctuations during the afternoon. Project 019-22 was completed and then we moved to project 056-22 to observe target W51D. Towards the end of the shift conditions were bad even for EMIR, so we stopped the observations shortly.

21:00-03:00; Karine & Isabelle: EMIR observations of 056-22 on the following sources: NGC1333, G188.94, G188.79, and Orion. High atmospheric fluctuations, with large tau variations 0.45 - 1.0.

Nov 15th, Tuesday

03:00-06:00; Philippe & Tuan-Anh: Observing with EMIR. Slight cloud cover, atmospheric opacity tau225 ~ 0.4-0.5, sky too unstable for NIKA2. Continued integrating on Orion for 056-22 until ~3:30 am (LT) as 079-22 targets were below 20 deg at the start of shift. Then switched to 079-22 at ~3:40 am (LT) and did 4 repetitions of onoff2 on C2Ea, 3 repetitions of onoff2 on C1a, 1 repetition of onoff on P07-C, and 1 repetition on P01-C.

06:00-10:30; Ioannis: EMIR observations for POLAMI LP (087-21).

DailyReportsNika2Pool081122 (last edited 2022-11-15 06:30:50 by NikaBolometer)